OpenStudy (anonymous):

Evaluate the following expression. 153^0

OpenStudy (alfie):

1

OpenStudy (anonymous):

Thank You (:

OpenStudy (alfie):

any number to the power of 0 is 0. Apart 0^0 which is an indeterminate form ;) you're welcome

OpenStudy (anonymous):

the answer is 1

OpenStudy (anonymous):

(-6)^-2 So This Is Something Like That, What Would That Be?

OpenStudy (anonymous):

.0277777778 You can round if necessary.

OpenStudy (anonymous):

or -24 if Im mistaken?

OpenStudy (anonymous):

Is the ^ used as a power of -2?

OpenStudy (alfie):

\[(-6)^{-2} = \frac{1}{-6^2} = \frac{1}{36}\]

OpenStudy (anonymous):

(: thank you. <3

OpenStudy (anonymous):

thanks